### v1

# Dual Computational Horizons: Incompleteness and Unpredictability in Intelligent Systems

Abstract
We formalize two independent computational limitations that constrain algorithmic intelligence: formal incompleteness and dynamical unpredictability. The former limits the deductive power of consistent reasoning systems while the later bounds long-term prediction under finite precision. We show that these two extrema together impose structural bounds on an agent’s ability to reason about its own predictive capabilities. In particular, an algorithmic agent cannot compute its own maximal prediction horizon generally. This perspective clarifies inherent trade-offs between reasoning, prediction, and self-analysis in intelligent systems.
keywords:
Algorithmic intelligence , Computability , Gödel incompleteness , Lyapunov exponent , Prediction horizon , Computational limits
journal: Information Processing Letters
1 Introduction
Modern AI has seen remarkable applications with groundbreaking results, yet fundamental theoretical limits remain underexplored. While scaling laws and architectures dominate research, intrinsic computational constraints from logic and dynamical systems place strict bounds on algorithmic intelligence. We identify and link two independent horizons:
  1. 1.
    Formal incompleteness: No consistent system capable of arithmetic can be deductively complete [godel, turing].
  2. 2.
    Dynamical unpredictability: Deterministic systems with positive Lyapunov exponents admit finite prediction horizons under finite-precision observations [ott, cohen2021].
We formalize their joint implications for algorithmic agents and connect them to phenomena such as multi-step reasoning limitations.
Related Work
Gödel incompleteness puts constraints on the formal reasoning in AI [godel, turing, bereska2024], while chaos theory establishes limits on long-term prediction [ott, cohen2021]. Computational irreducibility further restricts algorithmic shortcuts [wolfram]. Prior work often considers these separately; we explicitly link them in a unified computational framework. We do not claim new incompleteness results or new bounds in dynamical systems. We rather formalize their joint implications for algorithmic agents.
2 Formal Horizon: Limits of Deductive Reasoning
Let us take an example reasoning system, which can be represented as a pair $(\mathcal{S},\vdash)$ where $\mathcal{S}$ is a recursively enumerable theory over a countable language $\mathcal{L}$ and $\vdash$ denotes effective derivability. We assume $\mathcal{S}$ is consistent and capable of representing arithmetic.
Theorem 1 (Gödel’s First Incompleteness Theorem [godel]).
Let $\mathcal{S}$ be consistent and recursively enumerable. Then $\mathcal{S}$ is incomplete if there exists $\varphi\in\mathcal{L}$ such that neither $\mathcal{S}\vdash\varphi$ nor $\mathcal{S}\vdash\neg\varphi$.
This suggests that any agent whose reasoning is formalizable as a computable procedure cannot achieve deductive closure. In particular, such an agent cannot internally verify all truths expressible in its own language. This inherent limit is independent of the compute power of the agent.
3 Dynamical Horizon: Finite Prediction
Let $X$ be a metric space in $n$ dimensions, $X\subset\mathbb{R}^{n}$, and let
$x_{t+1}=F(x_{t}),$ (1)
be a deterministic dynamical system with maximal Lyapunov exponent $\lambda>0$. Suppose initial conditions are known only up to finite precision $\epsilon>0$, and let there be a computable predictor with an error $\epsilon$, given by $P_{\epsilon}$ such that:
$P_{\epsilon}:X\to X$ (2)
Let us consider $d$ as a distance metric in the phase space for the predicted and true states given by $P_{\epsilon}^{t}(x)$ and $F^{t}(x)$ respectively, at time $t$. Then, the prediction error grows at least as
$d(P_{\epsilon}^{t}(x),F^{t}(x))\geq C\epsilon e^{\lambda t}.$ (3)
For a fixed tolerance $\delta>0$, this lets us define the prediction horizon:
$T(\epsilon):=\frac{1}{\lambda}\log\frac{\delta}{C\epsilon}.$ (4)
Therefore, no algorithm can maintain arbitrary long-term prediction beyond $T(\epsilon)$.
4 Uncomputable Self-Prediction
Let us model an algorithmic agent [google2025, aws2025, ruffini2024] as $A=(R,P,M)$ with reasoning $R$, predictor $P$, and internal model $M$. Let
$T_{A}(\epsilon):=\text{maximal time $A$ can predict accurately.}$ (5)
Proposition 1 (Dual-Horizon Limitation).
Let $R$ be a consistent, recursively enumerable reasoning system. Then $R$ cannot compute $T_{A}(\epsilon)$ for arbitrary deterministic environments with $\lambda>0$.
Remark. The above argument is intended as a structural limitation rather than a complete computability-theoretic proof.
It is important to note here that neither incompleteness nor unpredictability due to chaos alone points towards Proposition 1. The limitation arises when an agent attempts to formally verify the bounds on its own predictive competence.
Proof Sketch. To compute $T_{A}(\epsilon)=\max\{t\mid\forall x\in X,\,d(F^{t}(x),P_{\epsilon}^{t}(x))<\delta\}$, the reasoner $R$ must decide, for every $t$, whether its predictor $P$ (parameterized by internal model $M$) satisfies the error bound over all initial states.
For hyperbolic systems ($\lambda>0$), this requires simulating $P^{t}(x)$ to arbitrary limits, but exponential error growth $d(F^{t}(x),P_{\epsilon}^{t}(x))\geq C\epsilon e^{\lambda t}$ implies no finite computation is sufficient for $t>T(\epsilon)$ [ott] for this simulation.
Moreover, $P$ and $M$ encode $R$’s own dynamics, making verification self-referential: deciding $P^{t}\approx F^{t}$ holds” is a $\Pi_{1}^{0}$ statement [enderton] in arithmetic. It is then equivalent to saying: $\forall$ (For all) trajectories, error remains bounded”. By Gödel’s second incompleteness theorem, consistent $R$ cannot prove its own consistency, hence cannot uniformly verify such simulation statements across all $t$.
Formally, the set $\{t\mid R\vdash(P^{t}\approx F^{t})\}$ is not computable, as its complement relates to the halting problem: $R$ cannot decide if divergent simulations “halt within $\delta$[turing]. Thus $R\not\vdash T_{A}(\epsilon)$ for arbitrary generality.
4.1 Illustrative Example: Self-Prediction of a Chaotic Simulator
Consider an agent $A$ whose environment model $M$ includes a deterministic chaotic map
$x_{t+1}=F(x_{t})$ (6)
with maximal Lyapunov exponent $\lambda>0$, implemented internally as part of its world simulator. Suppose $A$ attempts to compute its own prediction horizon $T_{A}(\epsilon)$ by simulating future states of $M$ using finite-precision arithmetic. Any such simulation necessarily introduces initial uncertainty $\epsilon>0$, causing prediction error to grow as $e^{\lambda t}$. Determining whether the simulation remains accurate beyond a candidate time $T$ requires $A$ to decide whether accumulated numerical error exceeds a fixed tolerance, a problem that reduces to undecidable self-verification when $M$ includes $A$ itself. Thus, algorithmic computation of $T_{A}(\epsilon)$ by $A$ is not possible universally.
Examples of such internally chaotic dynamics can be seen in continuous-depth neural networks and Neural Ordinary Differential Equation models, which are emerging as legitimate models to study neural networks. Consequently, these systems are not alien to the stability analysis in the light of dynamical systems either [haber2017stable].
5 Discussion
The dual-horizon perspective highlights a trade-off between stability of predictions and expressive/complex computational behavior. Systems operating near regimes of high sensitivity to initial conditions may achieve greater representational or optimization capacity, but at the cost of reduced long-term predictability. Conversely, constraining dynamics to improve predictability may limit expressivity.
From this viewpoint, phenomena such as error accumulation in long chains of reasoning may be understood as manifestations of finite prediction limits rather than implementation flaws. Modern neural language models, which operate as large-scale predictors under finite precision, provide an illustrative example: beyond certain depths, uncertainty accumulates and outputs may drift from intended linguistic outputs. We emphasize that this interpretation is illustrative rather than empirical.
Finally, incompleteness points towards inherent limits on self-verification. For sufficiently expressive reasoning systems, certain internal states cannot be conclusively classified using the system’s own rules. This aligns with recent analyses suggesting that absolute formal guarantees of safety or correctness may be unattainable for advanced autonomous agents, motivating probabilistic or external validation approaches [vassilev2025].
6 Conclusion
We identified two fundamental computational horizons, namely formal incompleteness and finite prediction under dynamical instability, and showed that their interaction imposes intrinsic limits on algorithmic intelligence. We present an argument that in general, no agent with computable reasoning and finite-precision prediction can determine its own maximal prediction limit. This result may act as a step towards the clarification on why long-horizon reasoning and self-verification inevitably degrade, even for highly capable systems. These limitations are structural rather than architectural and persist regardless of scale or implementation. We show that algorithmic intelligence does not overcome the bounds placed by incompleteness and chaos, but rather functions within these bounds. This approach provides us a principled, foundational basis for analyzing predictability, interpretability, and safety of modern algorithmically intelligent systems.